a coin is tossed 9 times what is the probability of getting ?
a)five heads
b)at least two heads

Answer by rothauserc(4718)

We use the binomial probability formula,
:
Probability (P) ( k successes in n trials ) = nCk * p^k * (1-p)^(n-k)
:
Note that nCk = n! / (k! * (n-k)!), p = 0.5
:
a) P ( k=5, n=9) = 9C5 * 0.5^5 * (1-0.5)^(9-5) = 0.246 approx 0.25
:
b) P ( at least 2 heads ) = 1 - P ( k=0, n=9) - P (k=1, n=9) = 1 - 0.002 - 0.020 = 0.978 approx 0.98
: